Logan Distributing Company of Atlanta sells fans and heaters to retail outlets throughout the Southeast. Joe Logan, the president of the company, is thinking about changing the firm's credit policy to attract customers away from competitors. The present policy calls for a 1/10, net 30 cash discount. The new policy would call for a 3/10, net 50 cash discount. Currently, 30 percent of Logan customers are taking the discount, and it is anticipated that this number would go up to 50 percent with the new discount policy. It is further anticipated that annual sales would increase from a level of $400,000 to $601,500 as a result of the change in the cash discount policy.

The increased sales would also affect the inventory level. The average inventory carried by Logan is based on a determination of an EOQ. Assume sales of fans and heaters increase from 14,810 to 22,550 units. The ordering cost for each order is $203, and the carrying cost per unit is $1.45 (these values will not change with the discount). The average inventory is based on EOQ/2. Each unit in inventory has an average cost of $12.

Cost of goods sold is equal to 65 percent of net sales; general and administrative expenses are 15 percent of net sales; and interest payments of 14 percent will only be necessary for the increase in the accounts receivable and inventory balances. Taxes will be 40 percent of before-tax income.

For average collection period, assume the customer pays on the last day possible (if they are getting the discount, that is day 10; if not, that is day 30 with the original policy and day 50 with the proposed policy).


  

Part A

Compute the accounts receivable balance before and after the change in the cash discount policy. Use the net sales (total sales minus cash discounts) to determine the average daily sales. Round your answer to the nearest whole dollar.

What is the accounts receivable balance before the change?
What is the accounts receivable balance after the change?    

Part B

Determine the EOQ before and after the change in the cash discount policy. Translate this into average inventory (in units and dollars) before and after the change in the cash discount policy. Round your answer to the nearest whole unit.

What was the EOQ prior to the change?
What is the EOQ after the change?        

What was the average inventory prior to the change?
What is the average inventory after the change?        

Part C

Determine the income after taxes before and after the new plan. Round your answer to the nearest whole dollar.

This income statement cannot be filled in online; it is provided as a possible approach to solving this problem.

Before Policy Change After Policy Change
Net sales (Sales - Cash discount)
Cost of goods sold (65%)
Gross Profit
Genreal and admin. expense (15%)
Operating profit
* Interest on increase in accounts receivable and inventory (14%)
Income before taxes
Taxes (40%)
Income after taxes

What was the income after taxes prior to the change?
What was the income after taxes after to the change?      

Part D

True or false: the firm should use the new cash discount policy.
Answer #1

Part (A)

We know, Accounts Receivable is equal to Average collection period * Average Daily Sales

Calculation of before the change: Average collection period = (0.30*10 days)+ (0.70*30 days)= 3+21 i.e. 24 days

Average daily sales= (Credit Sales- Discount)/360 days i.e. ($400000-$1200)/360 =$1107.78

Accounts Receivables before the change = 24 days * $1107.78= $26586.72

Calculation of Average Receivables after the change

Average Collection Period = (0.50*10 days)+ (0.50*50 days) i.e. 5+25 = 30 days

Average daily sales= (Credit Sales-discount)/360 days i.e. ($600000- $9000)/360 = $1641.67

Accounts Receivables after the change= 30 days* $1641.67= $49250.10 ( Average Collection Period* Average daily sales)

Part (B)

We know that EOQ= square root of (2SO)/C

EOQ prior to change = square root of (2*14810*203)/1.45 i.e. 6012860/1.45 = Square root of 4146800= 2036.36 units

EOQ after the change = square root of (2*22550*203)/1.45 i.e. 9155300/1.45 = square root of 6314000= 2512.76 units

Average Inventory is based on EOQ/2

Average Inventory before the change = 2036.36/2 i.e. 1018.18* $12 = $12218.16

Average Inventory after the change = 2512.76/2 i.e. 1256.38* $12 = $15076.56

Part (C)

Particulars Before Policy Change After Policy Change
Net Sales ( Sales- cash discount) $398800 $591000

Costs of goods sold (65%)

$259220 $384150
Gross Profit $139580 $206850
General and administration expense (15%) 59820 88650
Operating profit 79760 118200
Interest on increase in accounts receivable and inventory (14%) - 6031.27
Income before taxes 79760 112168.73
Taxes (40%) 31904 44867.49
Income after taxes 47856 67301.24

Working Notes: Calculation of Interest on Increase in Accounts Receivables and Inventory

Increase in Accounts Receivable = 14% of ($49250.10- $26586.72) i.e. 14% of $22663.38 = $3172.87

Increase in Inventory = 14% of ($15076.56- $12218.16) = $2858.4

Adding both the interest from Accounts Receivable and Inventory i.e. $3172.87 + $2858.4

Part (D) Yes firm should use the new cash discount policy as Income after taxes increases after the change of policy. Moreover increased operating profit after the policy change is higher than the interest charged on the increase inventory and accounts receivable. Therefore firm should adopt new cash discount policy.